OpenStudy (anonymous):

assume that $P is invested at 7% c.c. becomes $100 in a year USE A=Pe^rt a) Principal b) A(10)

OpenStudy (amistre64):

well, they give you A, r and t whats the holdup?

OpenStudy (anonymous):

well i just fixed it i have to find the A(10) for the second one but for the first equation i have 100/p=e^(.07)

OpenStudy (amistre64):

\[A=Pe^{rt}\] \[\frac{A}{e^{rt}}=P\]

OpenStudy (amistre64):

\[P = \frac{100}{e^{.07}}\] A(10) is when t=10 \[\Large A_{10} = \frac{100}{e^{.07}}e^{10*.07}\]

OpenStudy (anonymous):

ok thank you

OpenStudy (amistre64):

youre welcome

OpenStudy (amistre64):

just for practice :) \[\Large A_{10} = \frac{100}{e^{.07}}e^{10*.07}\] \[\Large A_{10} = \frac{100}{e^{.07}}e^{.70}\] \[\Large A_{10} = \frac{100}{e^{.07}}e^{.07+.63}\] \[\Large A_{10} = \frac{100}{\cancel{e^{.07}}}\cancel{e^{.07}}e^{.63}\] \[\Large A_{10} = 100~ e^{.63}\]

OpenStudy (anonymous):

oh ok i messed up on step lol thanks